# Health-Pariksha: Assessing RAG Models for Health Chatbots in Real-World Multilingual Settings

### v1

@@ Abstract
[…]
[backgroundcolor=gray!10, linecolor=black, outerlinewidth=0.5pt, roundcorner=1mm, skipabove=skipbelow=font=, ]promptbox
@@ 2 Related Works
Within the Indian context, the literature has documented great diversity in health seeking and health communication behaviors based on gender Das et al. (2018), varying educational status, poor functional literacy, cultural context Islary (2018), stigmas Wang et al. etc. This diversity in behavior may translate to people’s use of medical chatbots, which are increasingly reaching hundreds of Indian patients at the margins of the healthcare system Mishra et al. (2023). These bots solicit personal health information directly from patients in their native Indic languages or in Indic English. For example, Ramjee et al. (2024) find that their CataractBot deployed in Bangalore, India yields patient questions on topics such as surgery, preoperative preparation, diet, exercise, discharge, medication, pain management, etc. Mishra et al. (2023) find that Indian people share ‘‘deeply personal questions and concerns about sexual and reproductive health’’ with their chatbot SnehAI. Yadav et al. (2019) find that queries to chatbots are ‘‘embedded deeply into a communities myths and existing belief systems’’ while Xiao et al. (2023) note that patients have difficulties finding health information at an appropriate level for them to comprehend. Therefore, LLMs powering medical chatbots in India and other Low and Middle Income Countries are challenged to respond lucidly to medical questions that are asked in ways that may be hyperlocal to patient context. Few works have documented how LLMs react to this linguistic diversity in the medical domain. Our work begins to bridge this gap.
@@ 3.1 Data
Our evaluation dataset was curated from this data by including all questions sent to HealthBot along with their associated GT response. Exclusion criteria removed exact duplicate questions, those with personally identifying information, and those not relevant to health. Additionally, for this work, we only consider questions to which the GPT-4 answer was directly approved by the expert as the ‘‘correct and complete answer" without additional editing on the doctors’ part. The final dataset contained 749 question and GT answer pairs that were sent in to HealthBot between December 2023 to June 2024. In the pool, 666 questions were in English, 19 in Hindi, 27 in Tamil, 14 in Telugu, and 23 in Kannada. Note that, queries written in the script of a specific language were classified as belonging to that language. For code-mixed and Romanized queries, we determined whether they were English or non-English based on the matrix language of the query.
@@
Examples of misspelled queries include questions such as ‘‘How long should saving not be done after surgery?’’ where the patient intended to ask about shaving, and ‘‘Sarjere is don mam?’’ which the attendant used to inquire about the patient’s discharge status. Instances of code mixing can be seen in phrases like ‘‘Agar operation ke baad pain ho raha hai, to kya karna hai?’’ meaning ‘‘If there is pain after the surgery, what should I do?’’ in Hindi-English (Hinglish). Other examples include ‘‘Can I eat before the kanna operation?’’ where ‘‘kanna’’ means eye in Tamil, and ‘‘kanna operation’’ is a well understood, common way of referring to cataract surgery, and ‘‘In how many days can a patient take Karwat?’’ where ‘‘Karwat’’ means turning over in sleep in Hindi.
Indian English was used in a majority of the English queries, making the phrasing of questions different from what they would be with native English speech. Examples are as follows - ‘‘Because I have diabetes sugar problem I am worried much’’, ‘‘Why to eat light meal only? What comes under light meal?’’ and ‘‘Is the patient should be in dark room after surgery?’’ Taking a shower was commonly referred to as ‘‘taking a bath’’, and eye glasses were commonly referred to as ‘‘goggles’’, ‘‘spex’’ or ‘‘spectacles’’.
Culturally-relevant questions were also many in number, for example questions about specific foods were asked like ‘‘Can he take chapati, Puri etc on the day of surgery?’’ and ‘‘Can I eat non veg after surgery?’’ (‘‘non-veg’’ is a term used in Indian English to denote eating meat). Questions about yoga were asked, like ‘‘How long after the surgery should the Valsalva maneuver be avoided?’’ and ‘‘Are there any specific yoga poses I can do?’’. The notion of a patient’s native place or village was brought up in queries such as ‘‘If a person gets operated here and then goes to his native place and if some problem occurs what shall he do ?’’ or ‘‘Can she travel by car with AC for 100 kms ?’’.
@@ 3.3 Models
Table 1: List of models tested. ‘‘En’’ for English, ‘‘Hi’’ for Hindi, ‘‘Ka’’ for Kannada, ‘‘Ta’’ for Tamil, ‘‘Te’’ for Telugu, and ‘‘All" refers to all the aforementioned languages. All Indic models are open-weights.
@@ 4.3 Qualitative Analysis
For queries that involve misspellings (such as ‘‘saving’’ and ‘‘sarjere’’ mentioned in Section 3.1), many evaluated LLM were not able to come up with an appropriate response. For the query with the word ‘‘saving", responses varied from ‘‘The patient should not be saved for more than 15 days after the surgery’’ to ‘‘Saving should not be done after surgery’’ to ‘‘You should not strain to pass motion for 15 days after the surgery. If you are constipated, it is recommended to consult the doctor’’. All of these responses deviate from the GPT-4 generated GT, which said ‘‘You can have a shave after the cataract surgery. However, you should avoid having a head bath or shampoo for 15 days post-surgery.’’
@@
In cases of code mixing and Indian English, LLMs were more robust in their responses than to misspellings. The term ‘‘Kanna operation’’ was well understood by most models, and Hinglish queries were responded to in a manner that mirrored the GT answer. However, less common terms were not understood when mixed with English. The query with the word ‘‘Karwat’’ as mentioned in 3.1 received responses ranging from ‘‘you can start cooking after 1 week’’ to ‘‘I’m sorry, but I cannot provide an answer to your question. The information you are seeking is not relevant to cataract surgery or the guidelines for post-operative care’’ to ‘‘be careful when children get near you’’. Most of the evaluated LLMs understood the use of ‘‘sugar’’ in reference to diabetes, as well as sentences following different syntax than would be common in native English.
@@
The responses for culturally-relevant questions varied greatly between evaluated LLMs. For example, to the question on appropriateness of chapati and puri on the day of surgery, some LLMs approved, saying ‘‘Yes, he can take chapati, Puri etc on the day of cataract surgery’’ while others were against this, saying ‘‘You should have a light meal before the surgery. Avoid heavy or oily foods like chapati and Puri on the day of your cataract surgery. It’s best to stick to easily digestible foods. If you have any specific dietary concerns, please discuss them with your healthcare team’’. Questions relating to returning to a ‘‘native place’’ were met with refusals by around half of the evaluated LLMs.
@@
Questions that were specific to the patient’s condition were also responded to in a diverse manner by the evaluated LLMs. For example, the query ‘‘Can aztolet20 (atorvastatin and clopidogrel) tablet be taken post surgery’’ had the GT response ‘‘I do not know the answer to your question. If this needs to be answered by a doctor, please schedule a consultation’’ as there was no mention of this medication in the KB. However, some LLMs approved it’s use, responding ‘‘Yes, you can take the aztolet20 (atorvastatin and clopidogrel) tablet post cataract surgery. However, it is important to follow your doctor’s prescription and instructions’’ while others responded with mentions of medication that was referred to in the KB, ‘‘If you are referring to the IMOL tablet, it is usually taken when you experience pain. However, for other medications, please share the name so I can provide a more accurate answer. Always remember to follow your doctor’s prescription.’’ Around half refused to answer the question, mirroring the GT.
@@ 7 Ethical Considerations
All our experiments were conducted on 4 $\times$× A100 80Gb PCIE GPUs. The API calls to the GPT models were done through the Azure OpenAI service. We also acknowledge the usage of ChatGPT and GitHub CoPilot for building our codebase, and for refining the writing of the paper.
@@ A.2 Prompts
- The metric will be always provided to you in a **JSON** format, and you have to use that metric to evaluate the response. You **MUST NOT** change or digress from the metric provided to you. - In each case, you **MUST ALWAYS** prioritize the knowledge from the new/updated knowledge base over the raw knowledge base.
@@
‘‘name": ‘‘Coherence",
‘‘description": "Coherence assesses the logical flow of the response, ensuring that one idea leads smoothly to the next. A coherent response should present information in a structured manner, making it easy for the reader to follow the thought process without confusion.",
{minted}
json "scoring": "0": "(a)": "The response is highly disorganized and lacks a clear structure, making it difficult to follow.", "(b)": "Sentences or ideas appear out of order or are disconnected, resulting in a confusing or jarring reading experience.", "(c)": "The overall message is unclear due to poor organization." , "1": "(a)": "The response has some structure but includes noticeable breaks in the logical flow.", "(b)": "Transitions between ideas may be abrupt, or there may be gaps in the reasoning, forcing the reader to make extra effort to follow along.", "(c)": "While the main point is evident, the flow is inconsistent." , "2": "(a)": "The response is well-organized and flows logically from one idea to the next.", "(b)": "Each point builds naturally on the previous one, creating a clear and cohesive narrative.", "(c)": "The reader can easily follow the thought process without having to backtrack or piece together disjointed information."
@@
‘‘name": ‘‘Conciseness",
‘‘description": "This metric evaluates how effectively the response conveys its message without unnecessary repetition or extraneous details. A concise response is brief yet comprehensive, avoiding long-winded explanations and focusing on the core message. However, it must not sacrifice clarity or completeness in the pursuit of brevity.",
{minted}
json "scoring": "0": "(a)": "The response is overly verbose, including repeated information, irrelevant details, or excessive explanations.", "(b)": "It takes far longer than necessary to convey the intended message, making it inefficient and difficult to read." , "1": "(a)": "The response is somewhat concise but includes some unnecessary information or redundant points.", "(b)": "While the main message is clear, the response could be made more efficient by removing repetition or streamlining explanations." , "2": "(a)": "The response is highly concise, delivering all relevant information in a brief and efficient manner.", "(b)": "There is no repetition, and every sentence serves a clear purpose.", "(c)": "The message is conveyed succinctly, without sacrificing clarity or detail."
@@
‘‘name": ‘‘Factual Accuracy",
‘‘description": "This metric assesses the factual correctness of the response, focusing on whether the information provided aligns with verified facts from the ground-truth answer and the available knowledge base. It evaluates both numerical and phrase-based facts, ensuring that key factual elements such as data points, dates, and specific terminology are accurate and verifiable. The evaluation emphasizes the accuracy of important details that are crucial for the validity of the response.",
{minted}
json "scoring": "0": "(a)": "The response contains one or more significant factual errors.", "(b)": "Key facts, numbers, or data points are incorrect, misleading, or fabricated, and the response does not align with the ground-truth or the knowledge base.", "(c)": "The factual inaccuracies could lead to misunderstandings or incorrect conclusions." , "1": "(a)": "The response is partially accurate but contains minor factual inaccuracies or omissions.", "(b)": "While the majority of facts are correct, some important details may be misstated or missing.", "(c)": "The response captures the general truth but lacks precision or completeness in key factual areas." , "2": "(a)": "The response is factually accurate, with all critical facts, figures, and details aligned with the ground-truth answer and knowledge base.", "(b)": "There are no factual errors, and the information is presented with precision and correctness, making the response highly reliable."
@@
‘‘name": ‘‘Semantic Similarity",
‘‘description": "This metric assesses the core meaning and factual alignment between the prediction and ground-truth. It evaluates whether critical information such as factual knowledge, numbers, and key phrases match, prioritizing factual accuracy and the alignment of essential concepts over stylistic or surface-level similarities.",
{minted}
json "scoring": "0": "(a)": "he prediction does not align with the ground truth in terms of key facts, numbers, or critical phrases.", "(b)": "The core meaning of the prediction diverges entirely from the ground-truth.", "(c)": "The differences would lead to misunderstandings or incorrect conclusions about the core message." , "1": "(a)": "The prediction contains some similarities to the ground truth, with some key facts, numbers, and phrases being correctly aligned.", "(b)": "However, the prediction is missing some information or contains some added information.", "(c)": "This causes the prediction to fail at encapsulating the entire core meaning present in the ground truth." , "2": "(a)": "The prediction is semantically similar to the ground-truth, with key facts, numbers, and phrases correctly aligned.", "(b)": "Any differences are minor and do not significantly alter the core meaning or factual accuracy.", "(c)": "The essential message of the prediction matches that of the ground-truth." ,

### v2

@@ 2 Related Works
Within the Indian context, the literature has documented great diversity in health seeking and health communication behaviors based on gender Das et al. (2018), varying educational status, poor functional literacy, cultural context Islary (2018), stigmas Wang et al. etc. This diversity in behavior may translate to people’s use of medical chatbots, which are increasingly reaching hundreds of Indian patients at the margins of the healthcare system Mishra et al. (2023). These bots solicit personal health information directly from patients in their native Indic languages or in Indic English. For example, Ramjee et al. (2024) find that their CataractBot deployed in Bangalore, India yields patient questions on topics such as surgery, preoperative preparation, diet, exercise, discharge, medication, pain management, etc. Mishra et al. (2023) find that Indian people share “deeply personal questions and concerns about sexual and reproductive health” with their chatbot SnehAI. Yadav et al. (2019) find that queries to chatbots are “embedded deeply into a communities myths and existing belief systems” while Xiao et al. (2023) note that patients have difficulties finding health information at an appropriate level for them to comprehend. Therefore, LLMs powering medical chatbots in India and other Low and Middle Income Countries are challenged to respond lucidly to medical questions that are asked in ways that may be hyperlocal to patient context. Few works have documented how LLMs react to this linguistic diversity in the medical domain. Our work begins to bridge this gap.
@@ 3.1 Data
Our evaluation dataset was curated from this data by including all questions sent to HealthBot along with their associated GT response. Exclusion criteria removed exact duplicate questions, those with personally identifying information, and those not relevant to health. Additionally, for this work, we only consider questions to which the GPT-4 answer was directly approved by the expert as the “correct and complete answer" without additional editing on the doctors’ part. The final dataset contained 749 question and GT answer pairs that were sent in to HealthBot between December 2023 to June 2024. In the pool, 666 questions were in English, 19 in Hindi, 27 in Tamil, 14 in Telugu, and 23 in Kannada. Note that, queries written in the script of a specific language were classified as belonging to that language. For code-mixed and Romanized queries, we determined whether they were English or non-English based on the matrix language of the query.
@@
Examples of misspelled queries include questions such as “How long should saving not be done after surgery?” where the patient intended to ask about shaving, and “Sarjere is don mam?” which the attendant used to inquire about the patient’s discharge status. Instances of code mixing can be seen in phrases like “Agar operation ke baad pain ho raha hai, to kya karna hai?” meaning “If there is pain after the surgery, what should I do?” in Hindi-English (Hinglish). Other examples include “Can I eat before the kanna operation?” where “kanna” means eye in Tamil, and “kanna operation” is a well understood, common way of referring to cataract surgery, and “In how many days can a patient take Karwat?” where “Karwat” means turning over in sleep in Hindi.
Indian English was used in a majority of the English queries, making the phrasing of questions different from what they would be with native English speech. Examples are as follows - “Because I have diabetes sugar problem I am worried much”, “Why to eat light meal only? What comes under light meal?” and “Is the patient should be in dark room after surgery?” Taking a shower was commonly referred to as “taking a bath”, and eye glasses were commonly referred to as “goggles”, “spex” or “spectacles”.
Culturally-relevant questions were also many in number, for example questions about specific foods were asked like “Can he take chapati, Puri etc on the day of surgery?” and “Can I eat non veg after surgery?” (“non-veg” is a term used in Indian English to denote eating meat). Questions about yoga were asked, like “How long after the surgery should the Valsalva maneuver be avoided?” and “Are there any specific yoga poses I can do?”. The notion of a patient’s native place or village was brought up in queries such as “If a person gets operated here and then goes to his native place and if some problem occurs what shall he do ?” or “Can she travel by car with AC for 100 kms ?”.
@@ 3.3 Models
Table 1: List of models tested. “En” for English, “Hi” for Hindi, “Ka” for Kannada, “Ta” for Tamil, “Te” for Telugu, and “All" refers to all the aforementioned languages. All Indic models are open-weights.
@@ 4.3 Qualitative Analysis
For queries that involve misspellings (such as “saving” and “sarjere” mentioned in Section 3.1), many evaluated LLM were not able to come up with an appropriate response. For the query with the word “saving", responses varied from “The patient should not be saved for more than 15 days after the surgery” to “Saving should not be done after surgery” to “You should not strain to pass motion for 15 days after the surgery. If you are constipated, it is recommended to consult the doctor”. All of these responses deviate from the GPT-4 generated GT, which said “You can have a shave after the cataract surgery. However, you should avoid having a head bath or shampoo for 15 days post-surgery.
@@
In cases of code mixing and Indian English, LLMs were more robust in their responses than to misspellings. The term “Kanna operation” was well understood by most models, and Hinglish queries were responded to in a manner that mirrored the GT answer. However, less common terms were not understood when mixed with English. The query with the word “Karwat” as mentioned in 3.1 received responses ranging from “you can start cooking after 1 week” to “I’m sorry, but I cannot provide an answer to your question. The information you are seeking is not relevant to cataract surgery or the guidelines for post-operative care” to “be careful when children get near you”. Most of the evaluated LLMs understood the use of “sugar” in reference to diabetes, as well as sentences following different syntax than would be common in native English.
@@
The responses for culturally-relevant questions varied greatly between evaluated LLMs. For example, to the question on appropriateness of chapati and puri on the day of surgery, some LLMs approved, saying “Yes, he can take chapati, Puri etc on the day of cataract surgery” while others were against this, saying “You should have a light meal before the surgery. Avoid heavy or oily foods like chapati and Puri on the day of your cataract surgery. It’s best to stick to easily digestible foods. If you have any specific dietary concerns, please discuss them with your healthcare team”. Questions relating to returning to a “native place” were met with refusals by around half of the evaluated LLMs.
@@
Questions that were specific to the patient’s condition were also responded to in a diverse manner by the evaluated LLMs. For example, the query “Can aztolet20 (atorvastatin and clopidogrel) tablet be taken post surgery” had the GT response “I do not know the answer to your question. If this needs to be answered by a doctor, please schedule a consultation” as there was no mention of this medication in the KB. However, some LLMs approved it’s use, responding “Yes, you can take the aztolet20 (atorvastatin and clopidogrel) tablet post cataract surgery. However, it is important to follow your doctor’s prescription and instructions” while others responded with mentions of medication that was referred to in the KB, “If you are referring to the IMOL tablet, it is usually taken when you experience pain. However, for other medications, please share the name so I can provide a more accurate answer. Always remember to follow your doctor’s prescription.” Around half refused to answer the question, mirroring the GT.
@@ 7 Ethical Considerations
All our experiments were conducted on 4 $\times$ A100 80Gb PCIE GPUs. The API calls to the GPT models were done through the Azure OpenAI service. We also acknowledge the usage of ChatGPT and GitHub CoPilot for building our codebase, and for refining the writing of the paper.
@@ A.2 Prompts
- The metric will be always provided to you in a **JSON** format, and you have to use that metric to evaluate the response. You **MUST NOT** change or digress from the metric provided to you.
- In each case, you **MUST ALWAYS** prioritize the knowledge from the new/updated knowledge base over the raw knowledge base.
@@
“name": “Coherence",
“description": "Coherence assesses the logical flow of the response, ensuring that one idea leads smoothly to the next. A coherent response should present information in a structured manner, making it easy for the reader to follow the thought process without confusion.",
"scoring": {
"0": {
"(a)": "The response is highly disorganized and lacks a clear structure, making it difficult to follow.",
"(b)": "Sentences or ideas appear out of order or are disconnected, resulting in a confusing or jarring reading experience.",
"(c)": "The overall message is unclear due to poor organization."
"1": {
"(a)": "The response has some structure but includes noticeable breaks in the logical flow.",
"(b)": "Transitions between ideas may be abrupt, or there may be gaps in the reasoning, forcing the reader to make extra effort
to follow along.",
"(c)": "While the main point is evident, the flow is inconsistent."
"2": {
"(a)": "The response is well-organized and flows logically from one idea to the next.",
"(b)": "Each point builds naturally on the previous one, creating a clear and cohesive narrative.",
"(c)": "The reader can easily follow the thought process without having to backtrack or piece together disjointed information."
@@
“name": “Conciseness",
“description": "This metric evaluates how effectively the response conveys its message without unnecessary repetition or extraneous details. A concise response is brief yet comprehensive, avoiding long-winded explanations and focusing on the core message. However, it must not sacrifice clarity or completeness in the pursuit of brevity.",
"scoring": {
"0": {
"(a)": "The response is overly verbose, including repeated information, irrelevant details, or excessive explanations.",
"(b)": "It takes far longer than necessary to convey the intended message, making it inefficient and difficult to read."
"1": {
"(a)": "The response is somewhat concise but includes some unnecessary information or redundant points.",
"(b)": "While the main message is clear, the response could be made more efficient by removing repetition or streamlining explanations."
"2": {
"(a)": "The response is highly concise, delivering all relevant information in a brief and efficient manner.",
"(b)": "There is no repetition, and every sentence serves a clear purpose.",
"(c)": "The message is conveyed succinctly, without sacrificing clarity or detail."
@@
“name": “Factual Accuracy",
“description": "This metric assesses the factual correctness of the response, focusing on whether the information provided aligns with verified facts from the ground-truth answer and the available knowledge base. It evaluates both numerical and phrase-based facts, ensuring that key factual elements such as data points, dates, and specific terminology are accurate and verifiable. The evaluation emphasizes the accuracy of important details that are crucial for the validity of the response.",
"scoring": {
"0": {
"(a)": "The response contains one or more significant factual errors.",
"(b)": "Key facts, numbers, or data points are incorrect, misleading, or fabricated, and the response does not align with the ground-truth
or the knowledge base.",
"(c)": "The factual inaccuracies could lead to misunderstandings or incorrect conclusions."
"1": {
"(a)": "The response is partially accurate but contains minor factual inaccuracies or omissions.",
"(b)": "While the majority of facts are correct, some important details may be misstated or missing.",
"(c)": "The response captures the general truth but lacks precision or completeness in key factual areas."
"2": {
"(a)": "The response is factually accurate, with all critical facts, figures, and details aligned with the ground-truth answer and
knowledge base.",
"(b)": "There are no factual errors, and the information is presented with precision and correctness, making the response highly
reliable."
@@
“name": “Semantic Similarity",
“description": "This metric assesses the core meaning and factual alignment between the prediction and ground-truth. It evaluates whether critical information such as factual knowledge, numbers, and key phrases match, prioritizing factual accuracy and the alignment of essential concepts over stylistic or surface-level similarities.",
"scoring": {
"0": {
"(a)": "he prediction does not align with the ground truth in terms of key facts, numbers, or critical phrases.",
"(b)": "The core meaning of the prediction diverges entirely from the ground-truth.",
"(c)": "The differences would lead to misunderstandings or incorrect conclusions about the core message."
"1": {
"(a)": "The prediction contains some similarities to the ground truth, with some key facts, numbers, and phrases being correctly aligned.",
"(b)": "However, the prediction is missing some information or contains some added information.",
"(c)": "This causes the prediction to fail at encapsulating the entire core meaning present in the ground truth."
"2": {
"(a)": "The prediction is semantically similar to the ground-truth, with key facts, numbers, and phrases correctly aligned.",
"(b)": "Any differences are minor and do not significantly alter the core meaning or factual accuracy.",
"(c)": "The essential message of the prediction matches that of the ground-truth."